OpenStudy (anonymous):

Please help, I don't have a clue how to even begin solving this problem. The total U.S. consumption of energy in a particular year was 5.0*10^12 kilowatt-hours. What is the mass equivalence to this amount of energy?

OpenStudy (shamim):

5*10^12 kwh=5*10^12*1000*3600J Use E=mc^2 m=?

OpenStudy (shamim):

c=3*10^8m/s

OpenStudy (shamim):

E=energy m=mass

OpenStudy (shamim):

Response plz!!!

OpenStudy (anonymous):

So, if i am understanding this correctly, i need to divide my energy by c^2 to get my mass? And if so, what unit is my mass in?

OpenStudy (anonymous):

When i do the calculations the way that i think i need to do them, I end up with 200, my questions is, is that 200 grams or is it another unit?

OpenStudy (shamim):

Unit of m will b kg

OpenStudy (anonymous):

thanks for your help.

OpenStudy (shamim):

Medal??!!!